OpenStudy (anonymous):

\[F(x)=f(g(x)), \space f'(x)=\frac{x^2+1}{x-5}, \space g(x)=\sqrt{x}.\] Find F(x).

myininaya (myininaya):

so we need to find f so we need to integrate f' \[\int\limits_{}^{}\frac{x^2+1}{x-5} dx\]

myininaya (myininaya):

x+5 ---------- x-5| x^2+0x+! -(x^2-5x) ----------- 5x+1 -(5x-25) ---------- 26 so we have \[\int\limits_{}^{}(x+5+\frac{26}{x-5}) dx\]

OpenStudy (anonymous):

\[F =\rightarrow \sum_{^{_{_{\sqrt[\sqrt[?]{?}]{?}}^{?}}}}^{?}\]

myininaya (myininaya):

\[f(x)=\frac{x^2}{2}+5x+26 \ln |x-5|+C\] now we can find f(g(x))

OpenStudy (anonymous):

\[F(x)=\frac x2+5\sqrt{x}+26\ln |\sqrt{x}-5|+C\]

OpenStudy (anonymous):

looks good on the domain [0,25)u(25,infinity) as expected...

myininaya (myininaya):

Oh I didn't realize this was Broken Fixer's question. lol

OpenStudy (amistre64):

quick! delete it lol

myininaya (myininaya):

Broken Fixer knows everything.

OpenStudy (anonymous):

\[F(x)`=\frac 12+5\frac{1}{2\sqrt{x}}+\frac{13}{\sqrt{x}(\sqrt{x}-5)}\]